Jebediah:

How can I find out if an equation has no answer, has infinite answers has one answer or has 2 answers?

Jebediah:

@dude

Jebediah:

just a general question

dude:

Well if an equation has infinite solutions it is the same equation written in a different way 2(x+2)=2x+4 (infinite solutions) If 2 lines are parallel, or have the same coefficient they have no solutions \(x+7= x+4\)

Jebediah:

Oh okay

Jebediah:

So if they have 2 solutions, i have to go through looking for the second one?

dude:

Parabolas usually have 2 solutions https://www.desmos.com/calculator/5vm54e2hec

Jebediah:

Okay

Jebediah:

thanks

Jebediah:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Well if an equation has infinite solutions it is the same equation written in a different way 2(x+2)=2x+4 (infinite solutions) \(\color{#0cbb34}{\text{End of Quote}}\) But can it be like this? 2 + 3x = 2 + 3x?

Jebediah:

what if it is the same equation?

dude:

Yep, if it is the same equation it has infinite solutions

Jebediah:

oh okay

Jebediah:

thx

Jebediah:

:D

dude:

No problem :)

Jebediah:

wait

Jebediah:

so an equation like y = 5x + 10 y = 5x + 2 would this be no solution?

Jebediah:

sorry let me re-write that

Jebediah:

Equation A:y = 5x + 10 Equation B: y = 5x + 2

Jebediah:

@dude

dude:

Right

Jebediah:

so it would have no solution?

dude:

Correct

Jebediah:

ok

Jebediah:

wish i could give you 2 medals

Jebediah:

i have more questions probably XD

dude:

Hah thanks and okay